# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ V-B MIMO Fading Channels
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems [82, 38, 39, 78, 77, 37, 125, 76, 131], where, similar to the GMAC case, the idea is to divide the bit sequences into pilot and data parts. The transmitted signal of each user also consists of two parts (see Fig. 17 for details): the pilot part which is obtained by mapping the pilot bits to a non-orthogonal codebook [76, 37, 78, 77, 39, 125], or a codebook consisting of multiple stages of orthogonal pilots [38], and the data part which is the encoded by a a polar [38, 76, 37, 39] or an LDPC code [78, 77]. For decoding, as shown in Fig. 18, these schemes iteratively employ the following steps:
  • At the initial stage, they detect the active pilots using different algorithms such as generalized orthogonal matching pursuit (gOMP) [39, 125], AMP [76, 77, 131], energy detector [37], NP hypothesis testing [38] or CB-MLD [78].
  • Using the estimated channel coefficients, they obtain soft estimates of the transmitted codewords via MRC [76, 38, 39, 125], MMSE [37] or AMP [78], which are used for generating log-likelihood ratio (LLR) values to be fed a single-user decoder [38, 76, 37, 78, 39, 131] or a joint factor graph [77].
  • The contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the next iterations (this step is not employed by all the pilot-based schemes). Note that in [38, 37, 39, 125], the channel vectors are re-estimated using the successfully decoded codewords to be utilized for SIC.
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is extended to the fading scenario as well for the case of a single-antenna receiver in [111], and a massive MIMO receiver in [125]. The main idea is to allocate a part of the channel frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 19. Note that if the channel is a GMAC, the whole channel frame is used for data transmission.
@@
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently, hence the error from one block is propagated to the next one. To resolve this issue, there is also a new set of MIMO URA schemes that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal, and their corresponding channel coefficients [80, 79]. Specifically, in [79], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Besides, the authors of [80] propose to encode the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
@@
In Fig. 20, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients. The results in Fig. 20 show that due to adopting strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes (Fengler et al., 2022, Gkagkos et al., 2023, Ahmadi et al., 2024, Ozates et al., 2024) usually outperform the tensor-based (Decurninge et al., 2020) and CCS-based (Fengler et al., 2021) schemes for which sub-optimal coding schemes are utilized, while the SKP coding-based scheme in Han et al., 2023 following the Bayesian approach offers a similar performance to the best performing pilot-based schemes.
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution can be employing RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of the network by enabling line-of-sight paths between the transmitters and the receivers in problematic environments with many blocking obstacles [132, 133]. The authors in [134, 63] show that the use of RIS can improve user connectivity in URA by creating high-quality links between the BS and the users. More specifically, the RIS-based schemes in [134, 63] operate in two phases. In the first phase, called the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, called the data phase, transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [134] outperforms the one in [63], with a power saving of up to 10 dBm. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, and a more suitable metric in the RIS design in [134], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [63]. The authors improve the RIS design metrics of [134] in [135] and obtain some bounds on its performance.
@@
Conventionally, in MIMO URA systems, the ideal hardware units are assumed at both the BS and the UE side. However, due to the massive number of users and the antenna elements, the hardware needs to be inexpensive, and these elements are prone to HWIs such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWIs is considered in [136], where the solution developed in [39] is adapted to the presence of HWIs. Furthermore, a channel model with one single dominant path between a user and the BS is considered in [137], and an approach similar to the one in [131] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [138], namely, the users can start or end their transmission in any slot and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing.
Table IV: Summary of available coding schemes in URA with MIMO fading channels.
Paper Contributions Description
Gao et al., 2023 [87] Bounds Achievability and converse bounds.
@@
Fengler et al., 2021 [58] Coding scheme: CCS-based The pilots at each slot are found using a newly proposed coordinate-wise activity detector. Then, the detected pilots of different slots are stitched together using parity bits.
@@
Shyianov et al., 2021 [59] Coding scheme: CCS-based Instead of employing parity-check bits (as in [58]), the correlation between different slots is used for stitching pilots, hence, the energy efficiency is improved.
@@
Jiang et al., 2021 [129] Coding scheme: CCS-based A two-stage scheme with a long sub-message in first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage.
@@
Decurninge et al., 2020 [60] Coding scheme: tensor-based Users’ signal separation and detection is done by tensor decomposition.
@@
Luan et al., 2022 [62] Coding scheme: Tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [60].
Shao et al., 2022 [63] Coding scheme: tensor-based Users communicate with the BS through the RIS. For detecting the transmitted signals, a Bayesian approach is employed.
@@
Fengler et al., 2022 [76] Coding scheme: pilot-based A pilot and a polar codeword are transmitted sequentially. For decoding, the active pilots and corresponding channel coefficients are detected using MMV-AMP and MMSE, respectively, and a polar decoder is employed for data detection.
@@
Ahmadi et al., 2024 [38] Coding scheme: pilot-based Each user transmits its multiple stages of orthogonal pilots along with a polar codeword through a randomly selected slot. An energy detector, an MRC, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, an LS-based SIC removes the contribution of successfully decoded signals. Also, an efficient strategy is proposed for randomly dividing the users into different groups.
@@
Gkagkos et al., 2023 [37] Coding scheme: pilot-based A pilot and a random spread polar codeword are transmitted sequentially. An energy detector, an MMSE, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, a SIC removes the contribution of successfully decoded signals.
@@
Su et al., 2023 [77] Coding scheme: pilot-based A pilot is transmitted along with an LDPC coded and modulated codeword. At the receiver side, AMP is used for pilot detection and channel estimation, and an LDPC decoder is used for data detection. The selected pilot specifies the indexes over which every symbol of the LDPC codeword is transmitted.
@@
Su et al., 2023 [78] Coding scheme: pilot-based A pilot and an LDPC codeword are spread and transmitted consecutively. At the receiver’s end, an ML algorithm detects active pilots, an AMP technique estimates the corresponding channels, and a SIC block is applied following the LDPC decoder.
@@
Ozates et al., 2024 [39] Coding scheme: pilot-based The transmission frame is slotted, and each user randomly selects a slot to transmit a pilot along with a polar codeword. For decoding, a generalized OMP and an MMSE estimator are used for pilot detection and channel estimation, the polar decoder is used for data detection, and SIC removes the contribution of the decoded messages.
@@
Ozates et al., 2024 [125] Coding scheme: pilot-based The transmission frame is divided into pilot and data parts, and each user transmits a non-orthogonal pilot followed by its polar codeword based on ODMA. At the receiver, an iterative decoder employing gOMP, LMMSE, MRC, single-user decoding, and SIC is utilized.
Han et al., 2023 [79] Coding scheme: Bayesian approach Employs FEC code, and sparse spreading.
@@
Jiang et al., 2023 [80] Coding scheme: Bayesian approach Combines sparse regression codes for encoding and a message-passing algorithm for decoding.
@@
While most URA works consider the transmission over i.i.d. Rayleigh fading channels, the more practical Rician fading case is considered in [139] through an application of coded compressed sensing. A correlated fading scenario is tackled in [140]. Asynchronous MIMO URA is studied in [141, 143, 142]. A slotted concatenated coding scheme for the asynchronous URA is proposed in [141] based on the SPARCs and tree codes in an OFDM system. The authors of [142] tackle both the timing and frequency offsets by proposing an asynchronous URA scheme by employing SPARC, LDPC codes, and multistage orthogonal pilots. Moreover, symbol asynchronicity in URA is investigated in [143]. Moreover, some works have investigated URA in other communication setups such as cell-free MIMO [144] and erasure [145] channels. For instance, a cell-free setup is considered in [144] where the coverage area is divided into multiple regions each served by a receiver unit. The users in each region employ a common codebook different from other regions to transmit their messages, and the received signals at the receiver units are jointly processed in a central unit.
A summary of MIMO URA schemes is provided in Table IV.
@@ VI-A Asynchronous URA
In URA, the common assumption is that the users transmit simultaneously; or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of the users and devices in envisioned URA applications such as massive sensor networks, such transmission coordination is very challenging. This makes asynchronous URA an important research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. The reason is that due to the massive number of potential devices, and hence diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous.
There have been some attempts to address the asynchronous setup [73, 72, 116, 141, 142, 146, 93], most of which try to convert the problem to one similar to the synchronous case, e.g., by employing a limited decoding time-frame. For instance, in [73], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users. However, it is shown that if the asynchronicity is dealt with properly, it can lead to performance improvement even compared to the synchronous case [147]. One way to realize this is the use of frameworks that are more suitable for the asynchronous case. For instance, using sliding window frameworks with an (almost) infinite time horizon, instead of fixed-time time-frames.

### v2

@@ V-C3 Pilot-based schemes
V-C4 ODMA-based schemes
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is extended to the fading scenario as well for the case of a single-antenna receiver in [78], and a massive MIMO receiver in [110, 115, 116]. The main idea in [110] is to allocate a part of the time frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the time frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 16. At the receiver side, the gOMP algorithm is employed to detect the active pilots, followed by LMMSE channel estimation, MRC symbol estimation, single-user polar decoding, and SIC to recover the data part. Note that if the channel is a GMAC, the entire time frame is used for data transmission. On the other hand, the authors of [115] consider superimposed pilot and data transmission, where the pilot sequence occupies the entire time frame and data signal is transmitted by ODMA. The receiver operation starts with an energy detection-aided non-Bayesian pilot detection, and the same steps as in [110] are utilized for the decoding of the data part. A slotted ODMA-based scheme combining preamble transmission and convolutional LDPC coding is proposed in [117], with AMP decoding for the preamble and turbo probabilistic data association for multiuser detection of the data part at the receiver side. Furthermore, two slotted pilot-free ODMA schemes are proposed in [116] and [118]. While an iterative TIN strategy is employed for joint pattern and data detection in [116], random dictionary learning is studied in [118] in conjunction with clustering-aided activity detection.
V-C5 Bayesian approach
@@
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently. Hence, the error from one block is propagated to the next one. To resolve this issue, there is also a new set of MIMO URA schemes that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal, and their corresponding channel coefficients [120, 119]. Specifically, in [119], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Besides, the authors of [120] propose to encode the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
Figure 18: The decoding steps of the main URA approaches for MIMO fading channels.
Summary and Performance Comparison:
@@
A summary of MIMO URA schemes is provided in Table V, accompanied by the decoding steps of the main approaches summarized in Fig. 18. Moreover, in Fig. 19, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients. The results in Fig. 19 show that due to adopting strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes in [111, 12, 13, 14, 110] usually outperform the tensor-based scheme in [107] for which a sub-optimal coding scheme is utilized, while the SKP coding-based scheme in [119], following the Bayesian approach, offers similar performance to the best performing pilot-based schemes. Finally, note that, in order to benefit from increased spatial diversity, a receiver with a massive number of antennas (i.e., a massive MIMO setup) is considered in the provided numerical examples in most MIMO URA papers. However, in order for the suitable URA candidate for next-generation applications to cover a wider range of applications, a comprehensive comparison of the state-of-the-art schemes is required, especially for a smaller number of receiver antennas.
Computational Complexity:
@@
Table V: A summary of URA solutions for MIMO fading channels.
Paper Approach Description
Gao et al., 2023 [86] - Provides achievability and converse bounds.
@@
Fengler et al., 2021 [97] CCS-based The pilots at each slot are found using a newly proposed coordinate-wise activity detector. Then, the detected pilots of different slots are stitched together using parity bits.
@@
Shyianov et al., 2021 [98] CCS-based Instead of employing parity-check bits (as in [97]), the correlation between different slots is used for stitching pilots, hence, the energy efficiency is improved.
@@
Jiang et al., 2021 [104] CCS-based A two-stage scheme with a long sub-message in the first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage.
@@
Decurninge et al., 2020 [107] tensor-based Users’ signal separation and detection is done by tensor decomposition.
@@
Luan et al., 2022 [109] tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [107].
@@
Fengler et al., 2022 [111] pilot-based A pilot and a polar codeword are transmitted sequentially. For decoding, the active pilots and corresponding channel coefficients are detected using MMV-AMP and MMSE, respectively, and a polar decoder is employed for data detection.
@@
Ahmadi et al., 2024 [13] pilot-based Each user transmits its multiple stages of orthogonal pilots along with a polar codeword through a randomly selected slot. An energy detector, an MRC, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively, followed by an LS-based SIC.
@@
Gkagkos et al., 2023 [12] pilot-based A pilot and a random spread polar codeword are transmitted sequentially. An energy detector, an MMSE, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, a SIC removes the contribution of successfully decoded signals.
@@
Su et al., 2023 [112] pilot-based A pilot is transmitted along with an LDPC coded and modulated codeword. At the receiver side, AMP is used for pilot detection and channel estimation, and an LDPC decoder is used for data detection. The selected pilot specifies the indexes over which every symbol of the LDPC codeword is transmitted.
@@
Su et al., 2023 [113] pilot-based A pilot and an LDPC codeword are spread and transmitted consecutively. At the receiver’s end, a maximum likelihood detection (MLD) algorithm detects active pilots, an AMP technique estimates the corresponding channels, and a SIC block is applied following the LDPC decoder.
@@
Ozates et al., 2024 [14] pilot-based The transmission time frame is slotted, and each user randomly selects a slot to transmit a pilot along with a polar codeword. For decoding, a generalized OMP and an MMSE estimator are used for pilot detection and channel estimation, the polar decoder is used for data detection, followed by SIC.
@@
Ozates et al., 2024 [110] pilot-based The transmission time frame is divided into pilot and data parts, and each user transmits a non-orthogonal pilot followed by its polar codeword based on ODMA. At the receiver, an iterative decoder employing gOMP, LMMSE, MRC, single-user decoding, and SIC is utilized.
Han et al., 2023 [119] Bayesian Employs FEC code, and sparse spreading.
@@
Jiang et al., 2023 [120] Bayesian Combines sparse regression codes for encoding and a message-passing algorithm for decoding.
When comparing the computational complexity of the coding strategies for MIMO URA, CCS exhibits the lowest complexity among them, as its complexity is linear in $M_{r}$ and the number of common codebook elements [97]. Tensor-based [107] and non-slotted pilot-based [12] schemes have a similar complexity that is cubic in $K_{a}$. This complexity results from the canonical polyadic decomposition in the former and MMSE filtering for channel estimation in the latter. On the other hand, the complexity of slotted pilot-based schemes [13], [14] is cubic in the number of active users in the slot, which demonstrates the complexity advantage of slotting. Furthermore, the complexity of the Bayesian approach in [119] is linear in $K_{a}$, $M_{r}$, and $n$. To summarize, for practical system parameters, CCS has the lowest complexity, while the complexities of slotted pilot-based schemes and the Bayesian approach are similar and smaller than those of tensor-based and non-slotted pilot-based schemes.
V-D Existing URA Solutions on Alternative Setups
Most URA schemes in the literature consider direct links between all the users and the BS; however, in certain environments, the direct link between some users and the BS may be blocked or significantly attenuated. In these cases, one solution can be employing RIS, a promising technology that can furnish the URA system with high spectral efficiency and energy savings. Specifically, it is shown that a passive RIS equipped with many low-cost passive elements, which can intelligently tune the phase-shift of the incident electromagnetic waves, and reflect them in a desired direction without any amplification, can improve the efficiency of a URA system by enabling line-of-sight paths between the transmitters and the receivers in problematic environments with many blocking obstacles [121, 122]. More specifically, the RIS-based schemes in [123, 124] operate in two phases. In the first phase, called the RIS configuration phase, the BS detects the active pilots and estimates their respective channel coefficients, using which the BS suitably selects the phase shifts of the RIS elements. In the second phase, called the data phase, transmitted messages of active users are decoded using SIC-based decoders. The RIS-based URA solution in [123] outperforms the one in [124], with a power saving of up to 10 dBm. This improvement is due to the employment of a slotted transmission structure, which reduces the interference, as well as a more suitable metric in the RIS design in [123], where the overall phase-shift matrix is obtained by minimizing the decoding error probability, in contrast to maximizing the minimum channel gain among active users for the RIS design in [124]. The authors improve the RIS design metrics of [123] in [125] and obtain some bounds on its performance.
Conventionally, in MIMO URA systems, the ideal hardware units are assumed at both the BS and the UE side. However, due to the massive number of users and the antenna elements, the hardware needs to be inexpensive, and these elements are prone to hardware impairments (HWI) such as I/Q imbalance, phase noise, and quantization errors. The problem of URA with HWI is considered in [126], where the solution developed in [14] is adapted to the presence of HWI.
While most URA works consider the transmission over i.i.d. Rayleigh block-fading channels, some investigate other channel models. A channel model with one single dominant path between a user and the BS is considered in [127], and an approach similar to the one in [114] is proposed, exploiting the user location geometry by linking the transmission features such as the employed preamble or interleaver to the angle of arrival of the user. The dynamics of the user states and channels are taken into account in a slotted structure in [128], namely, the users can start or end their transmission in any slot and they can be active in a subset of the slots, and the channel coefficients change from one slot to another. The authors design a two-step decoding algorithm with an inner AMP and an outer variational message passing. Type-based MA, where each device observes the state of a process and transmits it to a common receiver is investigated in an unsourced setup in [129], where the receiver’s goal is to estimate the states and their multiplicity. The problem of user identification and authentication is applied to URA in [130] with the aid of a message authentication code. Moreover, the authors of [131] propose a URA scheme combining physical and medium access layer approaches where medium access layer slotting carries information. They show that carrying information by the medium access control layer improves the system performance. The Rician fading channel is considered in [132] through an application of coded compressed sensing. A correlated fading scenario is tackled in [133]. Frequency-selective fading is examined in the context of MIMO URA in [134] and [135]. The authors of [134] propose a tensor-based modulation scheme, while a slotted scheme employing OFDM is studied in [135].
The asynchronous setup receives increasing attention due to its easier implementation in hardware, and it is studied in MIMO URA in [136, 138, 137]. A slotted concatenated coding scheme is proposed in [136] based on the SPARCs and tree codes in an OFDM system. The authors of [137] tackle both the timing and frequency offsets by proposing an asynchronous URA scheme by employing SPARC, LDPC codes, and multistage orthogonal pilots. Moreover, symbol asynchronicity in URA is investigated in [138].
@@ V-D Existing URA Solutions on Alternative Setups
Furthermore, some works have investigated URA in other communication setups such as cell-free MIMO [139, 140, 141] and erasure [142, 143] channels. For instance, a cell-free setup is considered in [139] where the coverage area is divided into multiple regions each served by a receiver unit. The users in each region employ a common codebook different from other regions to transmit their messages, and the received signals at the receiver units are jointly processed in a central unit. Joint processing is also considered in [140] and [141], where a combination of channel coding and fixed sparse coding is employed in the former, while an uncoupled CCS scheme utilizing coarse location information for trellis stitching is proposed in the latter.
The common assumption in the URA literature is that all the users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, which is called multi-class URA is addressed in [144, 145] in the coded compressed sensing framework. More specifically, a scenario is considered where devices are clustered into two classes with different SNR levels and payload requirements. In the cluster with higher power, devices transmit using a two-layer superposition modulation, while those in the other cluster transmit with the same base constellation as in the first cluster. In order to recover packets at the receiver, a CCS-based approach is employed, where an outer code is used to stitch message fragments together across times and layers.
Now that we have reviewed the existing URA solutions, we provide a summary of different URA setups and the corresponding solution approaches in Fig. 20.
Figure 20: A summary of different URA setups in the literature and their corresponding common approaches.
@@ VI-A Asynchronous URA
In URA, the common assumption is that the users transmit simultaneously; or more precisely, their signals are received at the same time at the receiver (e.g., symbol synchronicity). However, due to the nature of the users and devices in envisioned URA applications such as massive sensor networks, such transmission coordination is very challenging. This makes asynchronous URA an important research direction to investigate. Note that even if the transmission synchronization is realized, the synchronous reception cannot be guaranteed. The reason is that due to the massive number of potential devices, and hence diverse distances to the receiver, the difference in reception times between many users can be larger than a time slot, which makes the setup asynchronous. It is shown that if the asynchronicity is dealt with properly, it can lead to performance improvement even compared to the synchronous case [146].
There have been some attempts to address the asynchronous setup [91, 90, 95, 136, 137, 51], however, with a bounded time horizon, which enable them to convert the problem to one similar to the synchronous case, e.g., by employing a limited decoding time frame. For instance, in [91], the authors employ a frequency-domain solution to convert the asynchronous problem to a phase-shift estimation one. Then, conventional synchronous strategies such as TIN-SIC are utilized to decode the active users.
The only work that focuses on a fully asynchronous (i.e., with an unbounded time horizon) URA setup is [147], where the authors consider a window-based decoding for GMAC URA. In this scheme, each user transmits a preamble (from a pool of available ones) for timing acquisition concatenated with a channel-coded payload, and the receiver uses a SIC-based sliding window for recovering the transmitted packets.
URA-specific challenges: The main URA-specific challenge of a fully asynchronous setup for fading URA is the computational complexity, which is much higher than the conventional (non-URA) setup due to the higher number of required preambles. More specifically, in an asynchronous GMAC setup, the preambles only serve to identify the start of the packets at the receiver, so a preamble pool of even a very small size would suffice; while in an asynchronous fading setup, they are also needed to estimate the channel coefficients, which are then used for user separation. In a conventional fading setup, the receiver only searches over the assigned preambles, whose number is equal to the number of active users. However, in a URA setup, due to the lack of coordination between the users and the receiver, the number of available preambles has to be much higher than in the conventional setup, hence much higher computational complexity for user activity detection and channel estimation. This computational complexity significantly increases in the asynchronous setup with a sliding window receiver, as the receiver shifts and correlates all possible preambles through the window to find the start of the packets as well as the user channel estimates, repeating this process with each shift of the sliding window. This highly increases the overall computational complexity, which can even make the solution practically infeasible. One way to reduce the complexity is to increase the moving steps of the sliding window, at the cost of lower overall performance. However, smarter schemes are needed to alleviate this issue.